# Learning to Equalize: Data-Driven Frequency-Domain Signal Recovery in Molecular Communications

### v2

@@ V Conclusion
In this letter, we proposed LSTM-FDE, a data-driven frequency-domain equalization scheme for MC systems based on LSTM. By leveraging LSTM’s ability to model temporal dependencies, the method achieves high detection accuracy without explicit channel modeling or estimation, learning directly from data. Simulation results show that LSTM-FDE outperforms both model-based FDE and other data-driven baselines across various symbol rates and under channel estimation errors. Compared with its time-domain counterpart, it provides a better trade-off between performance and complexity, making it suitable for high-rate MC scenarios. The approach also exhibits strong robustness in dynamic and noisy channels, highlighting its potential for practical biomedical applications such as in-body monitoring. Future work will explore lightweight recurrent architectures, MIMO extensions, and experimental validation on physical MC platforms.
Figure 6: The BER performance of LSTM-based TDE and FDE at different symbol durations.
@@ Appendix A LSTM-FDE Training and Inference Procedure
Experimental Environment: Intel i9-13900K CPU, NVIDIA GeForce RTX 4080 Laptop GPU, 128 GB RAM. Implemented in MATLAB 2023b. Dropout is applied to the recurrent layers for regularization.
1:Input: Training dataset $\mathcal{D}_{\text{train}}=\{(\mathbf{Y}_{i},b_{i})\}_{i=1}^{N}$,
2: Hyperparameters (Table II), $M=64$, $\epsilon=1\times 10^{-8}$, $S=10$
3:Output: Trained LSTM model with parameters $\mathbf{W}$
@@
6:   $\mathbf{y}(t)\leftarrow\text{time-domain received signal}$
7:   $\mathbf{Y}(m)\leftarrow\text{FFT}(\mathbf{y}(t),M)$ $\triangleright$ Frequency domain conversion
8:   $\widetilde{\mathbf{Y}}[m]\leftarrow\frac{|\mathbf{Y}[m]|-\text{Min}}{\text{Max}-\text{Min}+\epsilon}$ $\triangleright$ Min-max normalization to [0,1]
9:   Segment $\widetilde{\mathbf{Y}}$ into sequences of length $S$ $\triangleright$ Sliding window
@@
11:  Initialize LSTM (2 layers, 60 units, dropout=0.2) with random weights
12:  for epoch =1=1 to 3030 do
13:   for each batch of size 6464 do
@@
15:     $\mathcal{L}\leftarrow-\sum_{\ell}[b_{\ell}\log p_{\ell}+(1-b_{\ell})\log(1-p_{\ell})]$ $\triangleright$ BCE loss
16:     $\nabla_{\mathbf{W}}\mathcal{L}\leftarrow\text{BPTT}$ $\triangleright$ Backward pass
17:     $\mathbf{W}\leftarrow\mathbf{W}-\eta\cdot\text{Adam}(\nabla_{\mathbf{W}}\mathcal{L},\eta=0.001)$ $\triangleright$ Parameter update
@@
20:  return trained model
@@
22:procedure Inference($\mathbf{Y}_{\text{test}}$)
23:  Preprocess $\mathbf{Y}_{\text{test}}$ as in training steps 2–5
24:  $\mathbf{o}\leftarrow\text{LSTM}_{\mathbf{W}}(\widetilde{\mathbf{Y}}_{\text{test}})$ $\triangleright$ Forward pass (no dropout)
25:  $\hat{b}[m]\leftarrow\mathbf{1}_{o[m]\geq 0.5}$ $\triangleright$ Decision threshold
26:  return $\hat{b} \triangleright$ Decoded bit sequence

### v3

@@ V Conclusion
In this letter, we proposed LSTM-FDE, a data-driven FDE scheme for MC systems based on LSTM. By leveraging LSTM’s ability to model temporal dependencies, the method achieves high detection accuracy without explicit channel modeling or estimation, learning directly from data. Simulation results show that LSTM-FDE outperforms both model-based FDE and other data-driven baselines across various symbol rates and under channel estimation errors. Compared with its time-domain counterpart, it provides a better trade-off between performance and complexity, making it suitable for fast transmission scenarios. The approach also exhibits strong robustness in dynamic and noisy channels, highlighting its potential for practical biomedical applications such as in-body monitoring. Future work will explore lightweight recurrent architectures, MIMO extensions, and experimental validation on physical MC platforms.
Figure 6: BER performance comparison of LSTM-based TDE and FDE at different symbol durations.
The complete training and inference workflow of the proposed LSTM-FDE is summarized in the following pseudocode.
@@ Appendix A LSTM-FDE Training and Inference Procedure
1:Training dataset $\mathcal{D}_{\text{train}}=\{(\widetilde{\mathbf{Y}}_{i},b_{i})\}_{i=1}^{N}$
2:Hyperparameters (see Table II)
3:Trained LSTM model $\mathbf{W}$, Decoded bits $\hat{b}$
@@
6:   $y(t)\leftarrow\text{received signal according to (4)}$ $\triangleright$ Time-domain received signal
7:   $y[n]\leftarrow y(nT_{b}+t_{p})$ $\triangleright$ Sampling at peak time
8:   $\mathbf{Y}[m]\leftarrow\text{FFT}(\mathbf{y}[n])$ $\triangleright$ Frequency domain conversion
9:   $\widetilde{\mathbf{Y}}[m]\leftarrow\frac{|\mathbf{Y}[m]|-\text{min}}{\text{max}-\text{min}+\epsilon}$ $\triangleright$ Normalization according to (6)
@@
11:  Initialize LSTM with random weights $\mathbf{W}=\{\mathbf{W}_{f},\mathbf{W}_{i},\mathbf{W}_{c},\mathbf{W}_{o}\}$
12:  for epoch =1=1 to $E$ do
13:   for each batch $B$ do
@@
15:     $\mathcal{L}\leftarrow-\frac{1}{B}\sum_{\iota=1}^{B}\left[b_{\iota}\log p_{\iota}+(1-b_{\iota})\log(1-p_{\iota})\right]$ $\triangleright$ BCE loss according to (8)
16:     $\nabla_{\mathbf{W}}\mathcal{L}\leftarrow\text{BPTT}$ $\triangleright$ Backpropagation through time
17:     $\mathbf{W}\leftarrow\mathbf{W}-\eta\cdot\text{Adam}(\nabla_{\mathbf{W}}\mathcal{L})$ $\triangleright$ Parameter update
@@
20:  return trained model $\mathbf{W}$
@@
22:procedure Inference($\mathbf{y}_{\text{test}}[n]$)
23:  Preprocess $\mathbf{y}_{\text{test}}[n]$ as in Training $\triangleright$ FFT + normalization
24:  $\mathbf{o}\leftarrow\text{LSTM}_{\mathbf{W}}(\widetilde{\mathbf{Y}}_{\text{test}})$ $\triangleright$ Forward pass
25:  $\hat{b}[m]\leftarrow\mathbf{1}_{\{o[m]\geq 0.5\}}$ $\triangleright$ Threshold decision at 0.5
26:  return $\hat{b}$